COMPARISON OF PROFILES RETRIEVED FROM MEASUREMENTS BY THE ATMOSPHERIC CHEMISTRY EXPERIMENT (ACE) FOURIER TRANSFORM SPECTROMETER: TROPOSPHERE AND STRATOSPHERIC SPECIES

\underline{CURTIS P. RINSLAND}, NASA Langley Research Center Mail Stop 401A, Hampton, VA 23681 U.S.A. (c.p.rinsland@larc.nasa.gov); CHRIS D. BOONE, K. A. WALKER, B. F. BERNATH, Department of Chemistry, University of Waterloo, Waterloo, Ontario, Canada.

The Atmospheric Chemistry Experiment (ACE) was successfully launched in August 2003 to monitor the chemical composition Earth’s upper atmosphere. The primary instrument is a 0.02 cm^{−1} resolution Fourier transform spectrometer (FTS) measuring in the infrared from 750-4400 cm^{−1} with a maximum optical path difference of 25 cm. We describe preliminary comparisons of profile retrievals derived from spectra recorded in solar occultation mode based on algorithms developed at the University of Waterloo, Ontario, Canada, and at the NASA Langley Research Center, Hampton, VA U.S.A.